OpenStudy (anonymous):

A system of linear inequalities is shown below: x + y ≥ 4 y < 2x - 3 Describe the steps to graph the solution set to the system of inequalities. (10 points)

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

please help me

OpenStudy (dtan5457):

You would graph these two equations, and see where the shaded area stands for both

OpenStudy (dtan5457):

Have you ever done that?

OpenStudy (dtan5457):

the place where it's red and blue

OpenStudy (anonymous):

no i never done it

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

could you help me @Godlovesme

OpenStudy (anonymous):

u there

OpenStudy (godlovesme):

yes hang on i'll try

OpenStudy (anonymous):

ok thanks

OpenStudy (godlovesme):

np :)

OpenStudy (godlovesme):

y=mx+b u familiar with that (ik that's not an inequality :P)

OpenStudy (anonymous):

yes i am

OpenStudy (godlovesme):

so can u solve for y in the y+x >=4

OpenStudy (anonymous):

so u think that i could usey=mx+b

OpenStudy (godlovesme):

yep

OpenStudy (godlovesme):

 u got that? @123333333333333333333

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thats it

OpenStudy (godlovesme):

not yet

OpenStudy (godlovesme):

so when u graph inequalities step 1 - plot the y-intercept step 2- work on the slope

OpenStudy (godlovesme):

step 3- determine if the line is going to be solid or dashed step 4- shade above or below